OpenStudy (anonymous):

how do you solve y=-7/4x + 7/2. I need the slope, y intercept and x intercept. Then I have to graph it.

OpenStudy (anonymous):

The equation is in the form y=mx+b. In this form, m = the slope. b is the y-intercept. To find the x-intercept, set the equation equal to zero. y=-7/4x + 7/2 means m = -7/4, which is your slope 7/2 = b, which is your y-intercept. [y-intercept is (0,7/2)]. Set 0=-7/4x + 7/2 to find x-intercept, solving for x

OpenStudy (anonymous):

Hint for the x-intercept: It's not a whole number

OpenStudy (anonymous):

I am not sure how to do the fraction part of the y intercept.

OpenStudy (anonymous):

Okay, so the y-intercept is simply 7/2. The y-intercept is where x=0, and if you set x=0 in the equation you get y=7/2, which is your answer

OpenStudy (anonymous):

If it helps, don't think of it as a fraction, convert it into decimals: 7/2 = 3.5, so the y-intercept is 3.5

OpenStudy (anonymous):

Awesome thanks. I wasn't sure if I could just do that.